# Characterization of fractional Chern insulator quasiparticles in moiré transition metal dichalcogenides

### v1

@@ IV.1 Single delocalized quasiparticles
Using the energy levels in the presence of a single delocalized quasiparticle and the ground energy in the absence of quasiparticles, we can estimate the energy cost of a delocalized quasiparticle excitation. For a system of $N=N_{0},N_{s}=N_{s,0}=3N_{0}/2$ without quasiparticles, there are three nearly degenerate FCI ground states with mean energy ${\bar{E}}_{\rm gs}$. We define the creation energy of a single delocalized quasiparticle as $E_{\rm 1qp}({\bm{Q}})\equiv E_{N=N_{0}+1,N_{s}=N_{s,0}+1}({\bm{Q}})-{\bar{E}}_{\rm gs}$, where $E_{N=N_{0}+1,N_{s}=N_{s,0}+1}({\bm{Q}})$ is the energy level with momentum ${\bm{Q}}$ of the many-body Hamiltonian $H$ (impurity potential excluded) of the system size $N=N_{0}+1,N_{s}=N_{s,0}+1$ (such that one delocalized quasiparticle is added to the ground state). We refer to Ref. [89] for the allowed momentum in tilted samples. In our calculations, we do not include the contribution from a uniform neutralizing charge background, which was considered by previous studies of FQH states in Landau levels [91, 92, 93, 94]. In Fig. 5(a), we show $E_{\rm 1qp}({\bm{Q}})$ for $N_{0}=22$, whose ${\bar{E}}_{\rm gs}\approx 1.088\ {\rm eV}$. $E_{\rm 1qp}({\bm{Q}})$ can also be understood as a dispersive band of the single quasiparticle, with $\Delta_{s}$ the bandwidth. This quasiparticle band is quite flat at all twisted angles considered by us.
@@
where $\bm{Q}$ is no longer restricted to the discrete momenta imposed by the finite system size. Fig. 6(d) shows $E_{\text{1qp}}(\bm{Q})$ as a function of continuous $\bm{Q}$, which exactly reproduces the discrete values shown in Fig. 6(b). Interestingly, $E_{\text{1qp}}(\bm{Q})$ is periodic not only with respect to the moiré Brillouin zone defined by the underlying lattice, but also with respect to a smaller effective Brillouin zone, delineated by the dashed lines in Fig. 6(d). This emergent periodicity arises because $t(\bm{R})$ is dominated by hopping between sites separated by $3a_{M}$, leading to an effective Brillouin zone that is $3\times 3$ smaller than the original one. We note that this emergent periodicity in $E_{\text{1qp}}(\bm{Q})$ has recently been independently reported in Ref. [68]. Here we provide a perspective based on the anyon Wannier states and the hopping model.
We compare the interpolated $E_{\text{1qp}}(\bm{Q})$ in Fig. 6(d) to the values in Fig. 5(a) at the discrete momenta $\bm{Q}$ allowed in a different system size of $N=23,N_{s}=34$. The energy differences at each momentum can be as small as $3\times 10^{-4}$ meV and do not exceed $0.026$ meV. This close agreement validates the accuracy of the anyon hopping model in capturing the quasiparticle energy dispersion.
@@ IV.2 Two delocalized quasiparticles
Given the nearly flat single-quasiparticle band, the significant broadening of the subspace of two delocalized quasiparticles is a signal of the interaction between quasiparticles. For finite systems it is difficult to precisely subtract the contribution of single-quasiparticle dispersion from the two-quasiparticle spectrum, as their allowed momentum points are different. Here we only make a rough estimation of the interaction between two quasiparticles by calculating $V_{\rm 2qp}({\bm{Q}})\equiv E_{N=N_{0}+2,N_{s}=N_{s,0}+2}({\bm{Q}})-{\bar{E}}_{\rm gs}-2{\bar{E}}_{\rm 1qp}$, where $E_{N=N_{0}+2,N_{s}=N_{s,0}+2}({\bm{Q}})$ is the energy level with momentum ${\bm{Q}}$ of the many-body Hamiltonian of the system size $N=N_{0}+2,N_{s}=N_{s,0}+2$ (such that two delocalized quasiparticles are added to the ground state), and ${\bar{E}}_{\rm 1qp}$ is the $E_{\rm 1qp}({\bm{Q}})$ averaged over all momentum sectors. The result for $N_{0}=22$ is displayed in Fig. 5(b), whose ${\bar{E}}_{\rm 1qp}\approx 0.0884\ {\rm eV}$. Remarkably, $V_{\rm 2qp}$ is positive in all momentum sectors, with the lowest value $\approx 0.35\ {\rm meV}$. As the neglected single-quasiparticle bandwidth $\Delta_{s}$ is only $\approx 0.27\ {\rm meV}$ in this case, it is very likely that the two delocalized quasiparticles have a repulsive interaction.
The interactions between quasiparticles and quasiholes have been studied in the context of Landau levels [95, 96, 97, 98, 99]. We do similar analysis for the $\nu=2/3$ FQH quasiparticles in our LLL setup in Sec. II, using the same system sizes as in Fig. 5. In that case, we also find repulsive interactions between two quasiparticles. A recent work found the crossover from repulsive to attractive interaction between two $\nu=1/3$ Laughlin quasiholes when the range of the electron-electron interaction is decreased [99]. Our results in both the LLL and tMoTe2 are consistent with this observation, since the interaction is only weakly screened in our study.
@@
Figure 7: (a)-(c) Density distribution with a single-quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=25,N_{s}=37$. The impurity potential (indicated by the black cross) is located at (a) $\mathcal{R}_{M}^{M}$, (b) $\mathcal{R}_{M}^{X}$, and (c) $\mathcal{R}_{\rm Br}$, respectively. (d) Excess charges corresponding to the three cases in (a)-(c). We also give the data of the $\nu=2/3$ Coulomb FQH state [dashed line, the same as in Fig. 1(b)] for comparison. The horizontal reference line indicates $Q=e/3$, which is expected for a Laughlin-type quasiparticle.
@@ IV.3 Single localized quasiparticle
Let us first consider a single localized quasiparticle of the $\nu_{h}=2/3$ FCIs in tMoTe2. We are interested in the effect of this quasiparticle on particle’s real-space density, which is measured by $\Delta\rho({\bm{r}})=\rho({\bm{r}})-\rho_{0}({\bm{r}})$. Here $\rho({\bm{r}})$ is the particle’s density in the presence of a single quasiparticle, and $\rho_{0}({\bm{r}})$ is the particle’s density in the FCI ground state without quasiparticles [Fig. 3(a)]. In Figs. 7(a), 7(b), and 7(c), we show $\Delta\rho({\bm{r}})$ for $N=25,N_{s}=37$, which is the largest system size we can reach by ED. The impurity potential is located at $\mathcal{R}_{M}^{M}$, $\mathcal{R}_{M}^{X}$, and the bridge point $\mathcal{R}_{\rm Br}$ (middle of $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$), respectively. Extra particles indeed concentrate around the impurity potential in all of the three cases, which is a signal of a localized quasiparticle. With increasing distance from the impurity potential, $\Delta\rho({\bm{r}})$ oscillates and tends to zero. Unlike the FQH case [Fig. 1(a)], the density distribution in the presence of an FCI quasiparticle no longer has continuous rotational symmetry. Moreover, there is a clear dependence of $\Delta\rho({\bm{r}})$ on the position of the impurity potential. One can see that particles still prefer to occupy $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ regions as in the ground-state case, such that $\Delta\rho({\bm{r}})$ demonstrates $C_{6z}$ and $C_{3z}$ rotation symmetry, respectively, if the impurity potential is located at $\mathcal{R}_{M}^{M}$ and $\mathcal{R}_{M}^{X}$ sites [Figs. 7(a) and 7(b)]. The density profile for the impurity located at $\mathcal{R}_{M}^{M}$ closely follows that of the anyon Wannier state shown in Fig. 6(a).
@@
[…]
Figure 8: (a) The real-space Berry curvature probed by dragging a hole residing in the $K$-valley top valence band. (b) Braiding path (dashed line) of two quasiparticles for the system size $N=24,N_{s}=35$. The density distribution shown in the plot corresponds to two impurity potentials at the sample center and ${\bm{r}}=(0,0)$.
@@ IV.4 Braiding of two quasiparticles
We now examine the fractional statistics of quasiparticle excitations in the $\nu_{h}=2/3$ tMoTe2 moiré FCIs. We should emphasize that, while there is no physical magnetic field in moiré FCIs, a real-space Berry curvature $\Omega({\bm{r}})$ does exist, which contributes to the Berry phase when two localized quasiparticles are braided. This real-space Berry curvature has been reported for tight-binding models [69]. For tMoTe2, one can demonstrate $\Omega({\bm{r}})$ by mapping the tMoTe2 model to a Landau level problem, in which an effective periodic magnetic field with nonzero mean emerges [43]. To numerically probe this real-space Berry curvature, we adiabatically drag a single hole using a mobile impurity potential projected to the top valence band in the $K$ valley. If the hole’s wave function accumulates a Berry phase $\Phi({\bm{r}})$ along a closed path around point ${\bm{r}}$, we can approximate the real-space Berry curvature at ${\bm{r}}$ as $\Omega({\bm{r}})=\lim_{S({\bm{r}})\rightarrow 0}\Phi({\bm{r}})/S({\bm{r}})$, where $S({\bm{r}})$ is the area enclosed by the closed path. In Fig. 8(a), we show $\Omega({\bm{r}})$ for the $K$-valley top valence band of tMoTe2 at twist angle $\theta=3.7^{\circ}$. It is indeed very similar to the effective magnetic field obtained in Ref. [43]. For a single hole, the accumulated Berry phase along the boundary of a moiré unit cell is $2\pi$. Moreover, dragging a single localized quasiparticle along the same path gives Berry phase $32\pi/$ 3, which provides another compelling evidence of the fractional charge $e/3$ of the quasiparticle.
Then we fix one impurity potential at the sample center and move the other along a closed path, as shown in Fig. 8(b). We carefully choose the path such that the two impurity potentials are separated as far as possible and the path encloses an integer number of moiré unit cells. The technical method used to extract the Berry phase is the same as that in Sec. II.3. The accumulated Berry phase still contains two parts: the AB phase resulting from the real-space Berry curvature $\Omega({\bm{r}})$, and the anyon statistical phase. The AB phase is just $32\pi/$ 3 times the number of enclosed moiré unit cells. For the path in Fig. 8(b), we numerically find the anyon statistical phase as $1.311\pi$ for $N=24,N_{s}=35$, which has reasonably good agreement with the theoretical prediction $34\pi/$ 3. Again, the discrepancy is because the quasiparticles are not fully developed and separated in the finite system.
@@ V Trial Wave function for anyons in FCIs
where $z_{j}=x_{j}+iy_{j}$ is the complex coordinate of particle $j$, $N$ is the number of particles, and $\mathcal{N}_{L}$ is a normalization factor. Here we write the wave functions on the disk geometry for symbol simplicity. The quasihole excitation on top of the Laughlin state is given by
@@
where $g({\bm{k}})$ is the quantum metric of the band, and ${\rm tr}_{\tilde{g}}g\equiv\sum_{a,b=x,y}{\tilde{g}}_{ab}g^{ab}$ is the generalized trace with respect to a unimodular matrix $\tilde{g}$. For the top moiré valence band in tMoTe2, $T$ is approximately $0.1\sim 0.3$, as discussed in detail in Appendix D. In comparison, the LLL has exactly $T=0$ and momentum-independent quantum geometric tensor. More generally, an ideal Chern band with $T=0$ and momentum-dependent quantum geometries can be viewed as a generalized LLL with the wave functions of the form $\Theta_{\bm{k}}(\bm{r})=\mathcal{N}_{\bm{k}}\mathcal{B}(\bm{r})\Psi_{0,\bm{k}}(\bm{r})$, where $\bm{k}$ is the wave vector in the Brillouin zone, $\Psi_{0,\bm{k}}(\bm{r})$ is the magnetic Bloch wave function of the LLL, $\mathcal{B}(\bm{r})$ is a $\bm{k}$-independent but spatially-varying function, and $\mathcal{N}_{\bm{k}}$ is a normalization factor [79, 80, 81]. Since $T$ is small in tMoTe2, the Bloch wave function $\psi_{\bm{k}}(\bm{r})$ for its top moiré valence band in $K$ valley can be approximated by $\Theta_{\bm{k}}(\bm{r})$. We obtain the function $\mathcal{B}(\bm{r})$ by maximizing the momentum-averaged value of the overlap $|\langle\psi_{\bm{k}}|\Theta_{\bm{k}}\rangle|^{2}$ using a variational approach as outlined in Ref. [82]. The function $\mathcal{B}(\bm{r})$ captures the density fluctuations in the moiré band, and $|\mathcal{B}(\bm{r})|^{2}$ peaks at $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions similar to the density shown in Fig. 3(a).

### v3

@@ IV.1 Single delocalized quasiparticles
Using the energy levels in the presence of a single delocalized quasiparticle and the ground energy in the absence of quasiparticles, we can estimate the energy cost of a delocalized quasiparticle excitation. For a system of $N=N_{0},N_{s}=N_{s,0}=3N_{0}/2$ without quasiparticles, there are three nearly degenerate FCI ground states with mean energy ${\bar{E}}_{\rm gs}$. We define the creation energy of a single delocalized quasiparticle as
$E_{\rm 1qp}({\bm{Q}})\equiv E_{N=N_{0}+1,N_{s}=N_{s,0}+1}({\bm{Q}})-{\bar{E}}_{\rm gs},$ (23)
where $E_{N=N_{0}+1,N_{s}=N_{s,0}+1}({\bm{Q}})$ is the energy level with momentum $\bm{Q}$} of the many-body Hamiltonian $H$ (impurity potential excluded) of the system size $N=N_{0}+1,N_{s}=N_{s,0}+1$ (such that one delocalized quasiparticle is added to the ground state). We refer to Ref. [94] for the allowed momentum in tilted samples. In our calculations, we do not include the contribution from a uniform neutralizing charge background, which was considered by previous studies of FQH states in Landau levels [96, 97, 98, 99]. In Fig. 5(a), we show $E_{\rm 1qp}({\bm{Q}})$ for $N_{0}=22$, whose ${\bar{E}}_{\rm gs}\approx 1.088\ {\rm eV}$. $E_{\rm 1qp}({\bm{Q}})$ can also be understood as a dispersive band of the single quasiparticle, with $\Delta_{s}$ the bandwidth. This quasiparticle band is quite flat at all twisted angles considered by us. In usual Landau levels under a uniform magnetic field, the quasiparticle band is exactly flat. The weak dispersion observed here results from the non-uniform quantum geometry (Berry curvature and quantum metric) of the moiré band. Anyon dispersion in Landau levels under a non-uniform magnetic field (equivalent to non-uniform quantum geometry) was also reported recently [100].
@@
where $\bm{Q}$ is no longer restricted to the discrete momenta imposed by the finite system size. Fig. 6(d) shows $E_{\text{1qp}}(\bm{Q})$ as a function of continuous $\bm{Q}$, which exactly reproduces the discrete values shown in Fig. 6(b). Interestingly, $E_{\text{1qp}}(\bm{Q})$ is periodic not only with respect to the moiré Brillouin zone defined by the underlying lattice, but also with respect to a smaller effective Brillouin zone, delineated by the dashed lines in Fig. 6(d). This emergent periodicity arises because $t(\bm{R})$ is dominated by hopping between sites separated by $3a_{M}$, leading to an effective Brillouin zone that is $3\times 3$ smaller than the original one. We note that this emergent periodicity in $E_{\text{1qp}}(\bm{Q})$ has been theoretically anticipated in Ref. [101] and independently reported in numerical results of Ref. [72]. Here we provide a perspective based on the anyon Wannier states and the hopping model.
We compare the interpolated $E_{\text{1qp}}(\bm{Q})$ in Fig. 6(d) to the values in Fig. 5(a) at the discrete momenta $\bm{Q}$ allowed in a different system size of $N=23,N_{s}=34$. The energy differences at each momentum can be as small as $3\times 10^{-4}$ meV and do not exceed $0.026$ meV. This close agreement provides a benchmark for the accuracy of the anyon hopping model in capturing the quasiparticle energy dispersion, with hopping parameters truncated to $|\bm{R}|\leq 3a_{M}$.
@@ IV.2 Two delocalized quasiparticles
Given the nearly flat single-quasiparticle band, the significant broadening of the subspace of two delocalized quasiparticles is a signal of the interaction between quasiparticles. For finite systems it is difficult to precisely subtract the contribution of single-quasiparticle dispersion from the two-quasiparticle spectrum, as their allowed momentum points are different. Here we only make a rough estimation of the interaction between two quasiparticles by calculating
$V_{\rm 2qp}({\bm{Q}})\equiv E_{N=N_{0}+2,N_{s}=N_{s,0}+2}({\bm{Q}})-{\bar{E}}_{\rm gs}-2{\bar{E}}_{\rm 1qp},$
where $E_{N=N_{0}+2,N_{s}=N_{s,0}+2}({\bm{Q}})$ is the energy level with momentum $\bm{Q}$} of the many-body Hamiltonian of the system size $N=N_{0}+2,N_{s}=N_{s,0}+2$ (such that two delocalized quasiparticles are added to the ground state), and ${\bar{E}}_{\rm 1qp}$ is the $E_{\rm 1qp}({\bm{Q}})$ averaged over all momentum sectors. The result for $N_{0}=22$ is displayed in Fig. 5(b), whose ${\bar{E}}_{\rm 1qp}\approx 0.0884\ {\rm eV}$. Remarkably, $V_{\rm 2qp}$ is positive in all momentum sectors, with the lowest value $\approx 0.35\ {\rm meV}$. As the neglected single-quasiparticle bandwidth $\Delta_{s}$ is only $\approx 0.27\ {\rm meV}$ in this case, it is very likely that the two delocalized quasiparticles have a repulsive interaction.
The interactions between quasiparticles and quasiholes have been studied in the context of Landau levels [102, 103, 104, 105, 106]. We do similar analysis for the $\nu=2/3$ FQH quasiparticles in our LLL setup in Sec. II, using the same system sizes as in Fig. 5. In that case, we also find repulsive interactions between two quasiparticles. Motivated by a recent work which reported the crossover from repulsive to attractive interaction between two $\nu=1/3$ Laughlin quasiholes when the range of the electron-electron interaction is decreased [106], we have also examined much smaller $d$ ($d=2\ {\rm nm}$ for tMoTe2 and $d=\ell_{B}$ for the LLL), corresponding to stronger screening. Indeed, we find negative levels in $V_{\rm 2qp}({\bm{Q}})$, which is a signature of attractive interaction between quasiparticles [see Fig. 5(c) for the data in tMoTe2]. Therefore, our results in both the LLL and tMoTe2 are consistent with the observation in Ref. [106].
@@
Figure 7: (a)-(c) Density distribution with a localized single-quasiparticle excitation in the $\nu_{h}=2/3$ tMoTe2 FCI with $N=25,N_{s}=37$. The impurity potential (indicated by the black cross) is located at (a) $\mathcal{R}_{M}^{M}$, (b) $\mathcal{R}_{M}^{X}$, and (c) $\mathcal{R}_{\rm Br}$, respectively. (d) Excess charges corresponding to the three cases in (a)-(c). We also give the data of the $\nu=2/3$ Coulomb FQH state [dashed line, the same as in Fig. 1(b)] for comparison. The horizontal reference line indicates $Q=e/3$, which is expected for a Laughlin-type quasiparticle.
@@ IV.3 Single localized quasiparticle
Let us first consider a single localized quasiparticle of the $\nu_{h}=2/3$ FCIs in tMoTe2. Again we get three nearly degenerate ground states after diagonalizing the impurity potential in the subspace of a mobile quasiparticle. We are interested in the effect of this localized quasiparticle on particle’s real-space density, which is measured by
$\Delta\rho({\bm{r}})\equiv\rho({\bm{r}})-\rho_{0}({\bm{r}}).$ (28)
Here $\rho({\bm{r}})$ is the particle’s density in the presence of a single quasiparticle (averaged over the three localized quasiparticle states), and $\rho_{0}({\bm{r}})$ is the averaged particle’s density in the three FCI ground states without quasiparticles [Fig. 3(a)]. In Figs. 7(a), 7(b), and 7(c), we show $\Delta\rho({\bm{r}})$ for $N=25,N_{s}=37$, which is the largest system size we can reach by ED. The impurity potential is located at $\mathcal{R}_{M}^{M}$, $\mathcal{R}_{M}^{X}$, and the bridge point $\mathcal{R}_{\rm Br}$ (middle of $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$), respectively. Extra particles indeed concentrate around the impurity potential in all of the three cases, which is a signal of a localized quasiparticle. With increasing distance from the impurity potential, $\Delta\rho({\bm{r}})$ oscillates and tends to zero. Unlike the FQH case [Fig. 1(a)], the density distribution in the presence of an FCI quasiparticle no longer has continuous rotational symmetry. Moreover, there is a clear dependence of $\Delta\rho({\bm{r}})$ on the position of the impurity potential. One can see that particles still prefer to occupy $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ regions as in the ground-state case, such that $\Delta\rho({\bm{r}})$ demonstrates $C_{6z}$ and $C_{3z}$ rotation symmetry, respectively, if the impurity potential is located at $\mathcal{R}_{M}^{M}$ and $\mathcal{R}_{M}^{X}$ sites [Figs. 7(a) and 7(b)]. The density profile for the impurity located at $\mathcal{R}_{M}^{M}$ closely follows that of the anyon Wannier state shown in Fig. 6(a).
@@
as the length unit for tMoTe2. Under this choice, $d=10\ {\rm nm}$ corresponds to $d/\ell_{0}\approx 5$, which is consistent with the setting $d=5\ell_{B}$ in Sec. II. When $r\gtrsim 2\ell_{0}$, one can see oscillations of the excess charge around $e/3$ – the theoretical value of a Laughlin-type quasiparticle. This is a clear signal that the quasiparticle starts nucleating. However, $Q(r)$ does not converge even at the largest $r$ we can reach, suggesting that our system size is still insufficient for the quasiparticle to fully develop. The excess charge of $\nu_{h}=2/3$ moiré FCIs is similar to that of the $\nu=2/3$ FQH state when the length units are chosen as $\ell_{0}$ and $\ell_{B}$ for tMoTe2 and the LLL, respectively, but the oscillations of $Q(r)$ in the FQH case show quicker damping. If we take the radius of the $\nu=2/3$ FQH quasiparticle of Coulomb interaction as $6\ell_{B}$ as suggested by the braiding data in Sec. II.3, the $\nu_{h}=2/3$ FCI quasiparticle in tMoTe2 extents within a circle whose radius is at least $6\ell_{0}\approx 2.2a_{M}\approx 120\ \mathring$$\mathrm${A}}. This spatial extent could further increase if multiple bands are taken into account, as shown in Appendix D.
Figure 8: (a) The real-space Berry curvature probed by dragging a hole residing in the $K$-valley top valence band. (b) Braiding path (dashed line) of two quasiparticles for the system size $N=24,N_{s}=35$. The density distribution shown in the plot corresponds to two impurity potentials at the sample center and ${\bm{r}}=(0,0)$. We have averaged the density over the three nearly degenerate states with two localized quasiparticles.
@@ IV.4 Braiding of two quasiparticles
We now examine the fractional statistics of quasiparticle excitations in the $\nu_{h}=2/3$ tMoTe2 moiré FCIs. We should emphasize that, while there is no physical magnetic field in moiré FCIs, a real-space Berry curvature $\Omega({\bm{r}})$ does exist, which contributes to the Berry phase when two localized quasiparticles are braided. This real-space Berry curvature has been reported for tight-binding models [73]. For tMoTe2, one can demonstrate $\Omega({\bm{r}})$ by mapping the tMoTe2 model to a Landau level problem, in which an effective periodic magnetic field with nonzero mean emerges [43]. To numerically probe this real-space Berry curvature, we adiabatically drag a single hole using a mobile impurity potential projected to the top valence band in the $K$ valley. If the hole’s wave function accumulates a Berry phase $\Phi({\bm{r}})$ along a closed path around point $\bm{r}$}, we can approximate the real-space Berry curvature at $\bm{r}$} as
$\Omega({\bm{r}})=\lim_{S({\bm{r}})\rightarrow 0}\Phi({\bm{r}})/S({\bm{r}}),$ (30)
where $S({\bm{r}})$ is the area enclosed by the closed path. In Fig. 8(a), we show $\Omega({\bm{r}})$ for the $K$-valley top valence band of tMoTe2 at twist angle $\theta=3.7^{\circ}$. It is indeed very similar to the effective magnetic field obtained in Ref. [43]. For a single hole, the accumulated Berry phase along the boundary of a moiré unit cell is $2\pi$. Moreover, dragging a single localized quasiparticle along the same path gives Berry phase $2\pi/3$, which provides another compelling evidence of the fractional charge $e/3$ of the quasiparticle.
Then we fix one impurity potential at the sample center and move the other along a closed path, as shown in Fig. 8(b). We carefully choose the path such that the two impurity potentials are separated as far as possible and the path encloses an integer number of moiré unit cells. The technical method used to extract the Berry phase is the same as that in Sec. II.3. The accumulated Berry phase still contains two parts: the AB phase resulting from the real-space Berry curvature $\Omega({\bm{r}})$, and the anyon statistical phase. The AB phase is just $2\pi/3$ times the number of enclosed moiré unit cells. For the path in Fig. 8(b), we numerically find the anyon statistical phase as $1.311\pi$ for $N=24,N_{s}=35$, which has reasonably good agreement with the theoretical prediction $4\pi/3$. Again, the discrepancy is because the quasiparticles are not fully developed and separated in the finite system.
@@ V Trial Wave function for anyons in FCIs
where $z_{j}=x_{j}+iy_{j}$ is the complex coordinate of particle $j$, NN is the number of particles, and $\mathcal{N}_{L}$ is a normalization factor. Here we write the wave functions on the disk geometry for symbol simplicity. The quasihole excitation on top of the Laughlin state is given by
@@
where $g({\bm{k}})$ is the quantum metric of the band, and ${\rm tr}_{\tilde{g}}g\equiv\sum_{a,b=x,y}{\tilde{g}}_{ab}g^{ab}$ is the generalized trace with respect to a unimodular matrix $\tilde{g}$. For the top moiré valence band in tMoTe2, $T$ is approximately $0.1\sim 0.3$, as discussed in detail in Appendix D. In comparison, the LLL has exactly $T=0$ and momentum-independent quantum geometric tensor. More generally, an ideal Chern band with $T=0$ and momentum-dependent quantum geometries can be viewed as a generalized LLL with the wave functions of the form
$\Theta_{\bm{k}}(\bm{r})=\mathcal{N}_{\bm{k}}\mathcal{B}(\bm{r})\Psi_{0,\bm{k}}(\bm{r}),$ (34)
where $\bm{k}$ is the wave vector in the Brillouin zone, $\Psi_{0,\bm{k}}(\bm{r})$ is the magnetic Bloch wave function of the LLL, $\mathcal{B}(\bm{r})$ is a $\bm{k}$-independent but spatially-varying function, and $\mathcal{N}_{\bm{k}}$ is a normalization factor [83, 84, 85]. Since $T$ is small in tMoTe2, the Bloch wave function $\psi_{\bm{k}}(\bm{r})$ for its top moiré valence band in $K$ valley can be approximated by $\Theta_{\bm{k}}(\bm{r})$. We obtain the function $\mathcal{B}(\bm{r})$ by maximizing the momentum-averaged value of the overlap $|\langle\psi_{\bm{k}}|\Theta_{\bm{k}}\rangle|^{2}$ using a variational approach as outlined in Ref. [86]. The function $\mathcal{B}(\bm{r})$ captures the density fluctuations in the moiré band, and $|\mathcal{B}(\bm{r})|^{2}$ peaks at $\mathcal{R}_{M}^{X}$ and $\mathcal{R}_{X}^{M}$ positions similar to the density shown in Fig. 3(a).